OpenStudy (anonymous):

Nathan wrote the following steps to prove the equation: Given: 3(x + 2) = 6x + 3 Prove: x = 1

OpenStudy (anonymous):

Step Mathematical Statement Justification 0 3(x + 2) = 6x + 3 Given 1 3x + 6 = 6x + 3 Distributive Property 2 6 = 3x + 3 Commutative Property of Addition 3 3x + 3 = 6 Symmetric Property of Equality 4 3x = 3 Subtraction Property of Equality 5 x = 1 Division Property of Equality Which step of justification is incorrect, and what should the justification for that step be to solve the equation? Step 4; Commutative Property of Multiplication Step 3; Reflexive Property of Equality Step 1; Associative Property of Addition Step 2; Subtraction Property of Equality